OpenStudy (anonymous):

Strontium 90 is a radioactive material that decays according to the function A(t) = A0e ^-0.0244t where A0 is the initial amount present and A is the amount present at time t (in years) Assume that a scientist has a sample of 500 grams of strontium 90. What is the decay rate of strontium 90?

OpenStudy (unklerhaukus):

\[N(t)=N_0e^{-\lambda t}\] Where \(N(t)\) is the amount of \(N\) at time \(t\) , \(\lambda\) is the decay constant for the radioactive source \[A(t)=A_0e^{-0.2044t}\]\[A(0)=A_0e^{-0.2044\times 0}=A_0=500\text g\]

OpenStudy (unklerhaukus):

Is this the whole question? seams like some bits are missing,

OpenStudy (anonymous):

this is the who question

OpenStudy (anonymous):

whole

OpenStudy (unklerhaukus):

so you can find the decay rate/

OpenStudy (unklerhaukus):

radioactive decay is a first order reaction so the decay is at a constant rate

OpenStudy (anonymous):

isn't 500 the amount present?

OpenStudy (anonymous):

I think you have to use natural logs

OpenStudy (unklerhaukus):

Perhaps you are looking fot the Activity of the radioactive strontium

OpenStudy (anonymous):

i'm looking for the rate of decay

OpenStudy (unklerhaukus):

The activity is the number of decays per second ,so i guess you could call it the rate of decay

OpenStudy (anonymous):

ok.....and how do you get the rate of decay

OpenStudy (unklerhaukus):

\[\text{rate of decay}=\text{Activity}=\lambda N(t)\]

OpenStudy (anonymous):

that's not helping sorry

OpenStudy (unklerhaukus):

because \[\text{rate of decay}=\text{Activity}=-\frac{ \text dN(t)}{\text dt}\]

OpenStudy (anonymous):

well then how do you solve the problem. I'm not following

OpenStudy (unklerhaukus):

\[\text {activity}=-\frac{ \text dN(t)}{\text dt}=-\frac{ \text d}{\text dt}\left(N_0e^{-\lambda t}\right)\]

OpenStudy (unklerhaukus):

the decay rate means the rate of the number of decays , the negative of the amount of substance left after an amount of time

OpenStudy (unklerhaukus):

can you differentiate that ?

OpenStudy (anonymous):

it's not a differential problem it's a natural log problem

OpenStudy (unklerhaukus):

\[-\frac{ \text d}{\text dt}\left(N_0e^{-\lambda t}\right)=\lambda e^{-\lambda t}=\lambda N(t)\]

OpenStudy (unklerhaukus):

\[N(t)=N_0e^{-\lambda t}\]\[\frac{N(t)}{N_0}=e^{-\lambda t}\]\[\ln\left(\frac{N(t)}{N_0}\right)=\ln e^{-\lambda t}\]\[\ln\left(\frac{N(t)}{N_0}\right)=-\lambda t\]\[\ln\left(\frac{N_0}{N(t)}\right)=\lambda t\]

OpenStudy (unklerhaukus):

i hope i havent confused you with so many equations

OpenStudy (anonymous):

where you you put 500 grams in to the equation?

OpenStudy (unklerhaukus):

\(500 \text g\) is amount of strontium at \(t=0\) ie \[N(0)=N_0=500 \text g\]

OpenStudy (anonymous):

ok. and how do i solve the last step?

OpenStudy (unklerhaukus):

what are you up to

OpenStudy (unklerhaukus):

\[\text{activity}=\lambda N(t)=\lambda N_0e^{-\lambda t}\]\[=0.0244\left[\frac1{\text {yr}}\right]\times 500 [\text g]\times e^{-0.0244t}\]\[=\cdots\left[\frac{\text g}{\text{ yr}}\right]\]

OpenStudy (unklerhaukus):

the specific activity is \[\lambda N_0\]

OpenStudy (unklerhaukus):

which would just be \[=0.0244\left[\frac1{\text {yr}}\right]\times 500 [\text g]=\cdots \left[\frac {\text g}{\text {yr}}\right]\]

OpenStudy (unklerhaukus):

im not sure which the question is looking for.